solve for a log3 (2x+1)=4
log with lower three (2x+1)=4

Answer by nyc_function(2741)

log_3 (2x + 1) = 4
3^4 = 2x + 1
81 = 2x + 1
81 - 1 = 2x
80 = 2x
80/2 = x
40 = x